OpenStudy (anonymous):

Use the following equation: y = x2 The focus is located where?

OpenStudy (anonymous):

@SolomonZelman How do I know where the focus is located? Like how do I find that out

OpenStudy (bradely):

y = x^2 x^2=y 4a=1 a=1/4 Focus(0,a)=(0,1/4) Source: http://www.mathskey.com/question2answer/

OpenStudy (anonymous):

Wow thank you :D